OpenStudy (inkyvoyd):

for any positive natural number n that has a prime factorization without the numbers {2,5}is it possible to multiply it by another number such that the result is a string of nines? note that the number of nines does not matter, what matters is that the result is just nines.

OpenStudy (inkyvoyd):

for 1 - 1*9=9 for 3: 3*3=9 for 7: 7*7=49 for 9: 9*1=9

OpenStudy (inkyvoyd):

for 11: 11*9=99 for 13: 13*?=????

OpenStudy (inkyvoyd):

The issue with this question is how to find such a given number

OpenStudy (inkyvoyd):

*for 7: gimme a sec.

OpenStudy (inkyvoyd):

for 7: 7*142857=999999

OpenStudy (inkyvoyd):

Btw, I think I'll make this my problem of the day. It's simple math washed up a little so you don't recognize it.

OpenStudy (inkyvoyd):

@apoorvk